# Poisson Hail on a Wireless Ground

### v1

@@ 4.3 Reduction to a Lindley type recurrence relation
We have $D_{m}=T_{m}+W_{m}+\sigma_{m}$, where $\sigma_{m}$ is the time to serve all customers of block $m$ (or equivalently the time to complete the service time of the closing customer of this block) when its opening customer starts its service at time $T_{m}+W_{m}$. The key observation now follows from the construction by induction of Section 3, is that the random variable $\sigma_{m}$ is a measurable function of $W_{m}$ and $Y_{m}$:
@@
Note that $W_{m}$ is a function of $\{Y_{p}\}_{p<m}$ whereas conditionally on $W_{m}=W$, $\Sigma(W_{m},Y_{m})$ and $T_{m+1}-T_{m}$ are measurable functions of $Y_{m}$. Hence the dynamics of the sequence $\{W_{m}\}$ admits the representation
@@
$\widehat{\sigma}_{n}=\Sigma(\infty,Y_{n}).$
Note that the last random variable can be interpreted as follows: Think about all arrivals of block $Y_{n}$ to be already available at time 0. These customers then form a heap of hailstones. Start serving all customers of this heap following the induction steps of Section 3, namely in the order prescribed by the precedence relations between these customers and at any time, at the individual rate prescribed by the induction steps. Then $\widehat{\sigma}_{n}$ is the epoch at which the last element of this heap leaves.
Furthermore, since the random variables $\{Y_{n}\}$ are independent for $n\geq 1$ and identically distributed for $n\geq 2$, the same holds for $\widehat{\sigma}_{n}$: they are independent for $n\geq 1$ and identically distributed for $n\geq 2$.
@@
Under the foregoing assumptions, $\mathbb{E}\widehat{\sigma}_{m}<\infty,$ for any $m$. Moreover, there exists an absolute bound, say $K$, such that ${\mathbb{E}}\sigma_{m}(x)\leq K$, for all $x$ and $m$. In addition, the random variables $\{\sigma_{m}(x),m\geq 1,x\geq 0\}$ are uniformly integrable and, for any $m$, we have $\sigma_{m}(x)\to\widehat{\sigma}_{m}$ a.s., as $x\to\infty$. Therefore, for all $m$, ${\mathbb{E}}\sigma_{m}(x)\to{\mathbb{E}}\widehat{\sigma}_{1}$ as well.
@@
The number of customers in any block, say
$\nu_{m}:=n_{m}-n_{m-1}$
(see (7)), has a light-tailed distribution since it is nothing else than a geometrically distributed random variable with parameter
$p_{b}:={\mathbb{P}}(C_{n})=\left(\prod_{i=1}^{k}{\mathbb{P}}(B_{i}\subseteq S)\right)^{2}>0$
multiplied by $2k$. Further, since the total service rate (of all customers) in a non-empty system is bounded below by a positive constant, say $\kappa$ (see Remark 1), the total time for service of all customers in a block cannot exceed the sum of their heights, divided by $\kappa$. Note also that the random variables
@@
are i.i.d. Then, thanks to the Wald identity, we have that the mean total height of all customers in the block number $m=2,3,\ldots$ is equal to
${\mathbb{E}}\sum_{j=0}^{\nu_{m}-1}\sum_{r=0}^{2k-1}h_{n_{m-1}+2kj+r}$ $={\mathbb{E}}\sum_{j=1}^{\nu_{m}}\sum_{r=-k}^{k-1}h_{n_{m-1}+2kj+r}$
$={\mathbb{E}}\nu_{m}{\mathbb{E}}\sum_{r=0}^{2k-1}h_{r}=2k{\mathbb{E}}\nu_{1}{\mathbb{E}}h_{1}=:R_{1}<\infty.$
In the first block, only the distribution of the first $k$ random variables $\{h_{i}\}_{i=0}^{k-1}$ differs from that in the other blocks, in the sense that these random variables have the initial (unconditional) distribution. Then, clearly,
${\mathbb{E}}\sum_{j=0}^{\nu_{1}-1}\sum_{r=0}^{2k-1}h_{n_{m-1}+2kj+r}\leq R_{1}+k{\mathbb{E}}h_{1}=:R<\infty.$
Further, for any $m$ and $x$, the value of $\sigma_{m}(x)$ cannot exceed $D_{m}-D_{m-1}$ (recall that $D_{m}$ is the departure time of the last customer in block $m$) and that
@@
${\mathbb{E}}\sigma_{m}(x)\leq{\mathbb{E}}(T_{m}-T_{m-1})+R/\kappa=2k{\mathbb{E}}t_{1}+R/\kappa=:K<\infty.$
Next, the uniform integrability of all $\{\sigma_{m}(x)\}$ follows directly because they all are stochastically bounded by the sum of three random variables that are the same for all $\sigma_{m}(x)$ and that have finite expectations.
@@
$\sigma_{m}(x)=\widehat{\sigma}_{m},\ \ \text{ for all}\ \ x>T_{m}-T_{m-1}.$
Indeed, $T_{m-1}+x$ is the time when the last customer in block $m-1$ leaves the system, so if $x>T_{m}-T_{m-1}$, then all customers from block $m$ are already in the system at time $T_{m-1}+x$, like in the saturated system. ∎
4.4 Main Stability Results
We obtain two stability results, one for the Markov chain $\{W_{n}\}$ and one for the continuous-time process $\{W(t),t\geq 0\}$ admitting $\{W_{n}\}$ as an embedded Markov chain.
Below, we use the following notation: $T=T_{1}$, $\nu=\nu_{1}$ (where the latter is defined in the proof of Lemma 10), and $\widehat{\sigma}=\widehat{\sigma}_{1}$. Since $T=\sum_{i=1}^{\nu}t_{i}$, ${\mathbb{E}}\nu=2k/p_{b}$, $\lambda=1/(|{\cal X}|\mathbb{E}t_{1})$ and $\nu$ does not depend on $\{t_{i}\}$, we get
${\mathbb{E}}T=\frac{2k}{\lambda|{\cal X}|p_{b}}.$ (11)
@@ 4.4 Main Stability Results
    If ${\mathbb{E}}T>{\mathbb{E}}\widehat{\sigma}$ – or, equivalently,
    $\lambda<\lambda_{c}:=\frac{2k}{p_{b}|{\cal X}|{\mathbb{E}}\widehat{\sigma}},$ (12)
@@
    In addition, if the distribution of $t_{1}$ has infinite support, then the Markov chain $\{W_{n}\}$ admits a single stationary distribution, say $\pi$, and, for any initial state $x\in{\cal X}$, the distribution of the Markov chain converges to $\pi$ in the total variation norm.
@@
    If ${\mathbb{E}}T<{\mathbb{E}}\widehat{\sigma}$ and if there is a constant $\gamma>0$ such that, for any $x$
    ${\mathbb{P}}_{\psi}^{0}(\Sigma(x,Y_{1})>T_{1}+\gamma)\geq\gamma,$ (13)
@@
[…]
@@
In some special cases, there are closed-form expressions for $p_{b}$ and ${\mathbb{E}}\widehat{\sigma}$. For example, for $S_{i}=b(x_{i},R_{i})$ where $\{R_{i}\}$ are exponentially distributed i.i.d. with mean radius $R$, it is sufficient to define the zigzag event based on a single arrival for which the exclusion set covers $\cal X$, that is, $C_{n}=\{{\cal X}\subset S_{n}\}$. In this case, $p_{b}$ is simply the tail probability of an exponential random variable, and we have:
$\lambda_{c}=\frac{\exp(\sqrt{2}W/R)}{|\cal X|{\mathbb{E}}\widehat{\sigma}}.$
The tractability of ${\mathbb{E}}\widehat{\sigma}$ depends both on the distribution of $S_{i}$ and the pathloss function $\ell$. In general, we can estimate $p_{b}$ and ${\mathbb{E}}\widehat{\sigma}$ by Monte Carlo simulation. The details are given in Section 5.
For any time $t\geq 0$, consider an auxiliary system in which we stop arrivals from time $t$ onward and admit only customers arriving before time $t-0$. In this system, there is a (random) finite number of arrivals. Denote by $Z(t)$ the time when the last customer leaves such a system and by $W(t)=Z(t)-t$ the time to empty the system after the interruption of the input process. One can see that $Z(t)$ is finite a.s. (in particular, because the sum of the service rates of customers on the ground is bounded from below) and that $W_{n}=W(T_{n})$, for $n=1,2,\ldots$.
@@
    Assume the assumptions of item (i) of the previous theorem hold. If in addition the distribution of $t_{1}$ has a non-lattice distribution, then the Markov process $W(t)$ admits a single stationary distribution and, for any initial finite configuration of customers, the distribution of $W(t)$ converges weakly to its stationary version. If in addition the distribution of $t_{1}$ is spread-out, the convergence holds in the total variation norm.
@@
[…]
The proofs are given in Appendix 8.
@@
5 Simulations
5.1 Estimates of the stability threshold
@@
In view of the expression obtained in (12), the most natural way to estimate $\lambda_{c}$ is to estimate $p_{b}$ and $\mathbb{E}\widehat{\sigma}_{1}$ by Monte Carlo simulation. Both problems can be solved by producing perfect block samples and using them as input to the induction steps of Section 3. For this purpose, in these blocks, there is no need to keep the arrival epochs. The simplest way is to first sample an i.i.d. sequence of arrival loci, height, and exclusion sets until one sees a first $C$ event and to proceed sampling such data until one sees a second $C$ event. This provides of sample of a block by restriction to the arrivals that start with the zag of the first $C$ event and lasts until the zig of the second one. This provides a sample of the number of arrivals in a block, say $N$. One then takes this block as input to the induction steps of Section 3. This in turn provides a perfect sample of $\hat{\sigma}_{1}$. By repeating this, one gets an estimate (together with confidence intervals) of $\mathbb{E}N=1/{p_{b}}$ and of $\mathbb{E}\widehat{\sigma}$.
5.2 Examples
@@
This subsection defines two sets of basic examples, one in the discrete space setting and the other in the continuous space setting.
5.2.1 Discrete-Space Setting
We define here an example in the 1D discrete-space setting. Let $N$ be an even natural number $N\geq 4$. There are $N$ servers located equidistantly on the circle of length $N$. We number them $1$ to $N$ clockwise, say. Customers are represented as rectangles $M\times h$ centered at $x$. Here, $x$ represents the “center” of the customer, which is one of the locations of the servers. Here, $M$ is a fixed odd integer, with $1\leq M\leq N$. The meaning of $M$ is that the customer needs simultaneous service of $M$ neighboring servers. More precisely, the exclusion set of a customer centered at $x$ concerns the $(M-1)/2$ servers on the left and the $(M-1)/2$ servers on the right of server $x$. Exclusion means that if two customers have common servers, then they cannot be served simultaneously. As in the previous discussions, we assume here that customers are served in a locally FCFS way. Finally, $h$ is a positive random variable which represents the file size of a customer.
The distance between two customer centers is the minimum of the two distances between them, clockwise and counterclockwise. Let $d$ be a fixed positive integer. The attenuation function at distance $u\in\mathbb{N}$ is $l(u)=v\cdot 1_{u\leq d}$, where $v>0$ is a constant. The thermal noise is equal to 1. Hence, the service rate of a customer at $x$ and in service is equal to $\frac{1}{1+v\phi_{d}(x)}$ with $\phi_{d}(x)$ with $\phi_{d}(x)$ the number of customers being served at a distance less than or equal to $d$, without counting the customer at $x$.
Note that for $(M-1)/2\geq d$, this model coincides with Poisson Hail since the rate of any customer being served is always 1. It is known that Poisson Hail is strictly monotonic in the footprint of the hailstones. More precisely, for $N$ and $d$ fixed, the value of $\lambda_{c}$ is a decreasing function of $M$, in the range $M\geq 2d+1$.
A special case of this model is obtained by taking $d=N/2$. That if a customer’s service “interferes” with at least another customer, then it interferes with exactly one customer, and its service rate is then $\frac{1}{v+1}$.
If $M>d$, then there is no way to serve two or more customers simultaneously, and we have a single-server queue with service rate 1. If $M=d$, then the simultaneous service of two customers may only occur if the distance between the customer centers is exactly $d$, and at most two customers are served simultaneously.
Consider the saturated system (it means that there are infinitely many customers in the queue in front of the system, which are numbered $1,2,\ldots$) in the case $M=d$. We assume that customer $i$ is centered at $x_{i}$ and brings $H_{i}$ units of work to serve, that $\{x_{i}\}$ is an i.i.d. sequence with elements uniformly distributed on the discrete set of $N$ locations, and that $\{h_{i}\}$ is an i.i.d. sequence of positive random variables.
5.2.2 Continuous-Space Setting
Consider the PHWG model defined on the continuum in 1D. The state space is an interval ${\cal X}=[-W,W]$ of fixed length $2W$. Arrivals follow a space-time Poisson process on ${\cal X}$ with intensity $\lambda$. The ends are glued together to form a circle. In 2D, we consider the state space as a square window ${\cal X}=[-W,W]^{2}$. Arrivals follow a space-time Poisson process on $\cal X$ with intensity $\lambda$. Similarly, the torus version of the model is used to evaluate the wireless scenario.

### v2

@@ III-C Reduction to a Lindley-Type Recurrence Relation
We have $D_{m}=T_{m}+W_{m}+\sigma_{m}$, where $\sigma_{m}$ is the time to serve all customers of block $m$ (or equivalently the time to complete the service time of the closing customer of this block) when its opening customer starts its service at time $T_{m}+W_{m}$. The key observation now follows from the construction by induction of Section II that the random variable $\sigma_{m}$ is a measurable function of $W_{m}$ and $Y_{m}$:
@@
It should be clear that $\sigma_{m}$ and the dynamics of block $m$ depend on the waiting time $W_{m}$. To see this, consider two subsequent arrivals in this block whose exclusion sets do not intersect with each other. If $W_{m}$ is sufficiently large (e.g., $\infty$), then at the beginning of the service (end of waiting), these two arrivals will be served at the same time. If $W_{m}$ is small (e.g., 0), on the other hand, then the two arrivals may be served one by one, each enjoying the maximum transmission rate free of interference.
Fig. 3 illustrates the block structure based on a deterministic set of $\{B_{i}\}_{i=1}^{k}$ in $\cal X\subset\mathbb{R}$ for $k=4$. In Fig. 3 (a), the base of a rectangle indicates the arrival time and the height indicates the file size. There is one (non-zigzag) set of arrivals of size $2k$ between two zigzags, indexed by their order of arrival. Arrivals with the same color have the same exclusion radius and file size. In Fig. 3 (b), for $W_{m}\geq\widehat{T}_{m}-T_{m}$, all arrivals are already present at the beginning of the service of the block, hence the block is maximally compressed while preserving the local FCFS mechanism.
@@
Note that $W_{m}$ is a function of $\{Y_{p}\}_{p<m}$, whereas, conditionally on $W_{m}=W$, $\Sigma(W_{m},Y_{m})$ and $T_{m+1}-T_{m}$ are measurable functions of $Y_{m}$. Therefore, the dynamics of the sequence $\{W_{m}\}$ admits the representation
@@
$\quad\Sigma_{m}(x)=\Sigma(x,Y_{m})\quad\text{and}\quad\widehat{\sigma}_{m}=\Sigma(\infty,Y_{m}).$
Two observations are in order:
  • Each of the maps $x\to\Sigma_{m}(x)$ from $\mathbb{R}^{+}$ to $\mathbb{R}^{+}$ can be built using the construction of the last section as a piecewise affine random stochastic process. In particular, the last map is measurable.
  • The family of stochastic processes $\{\Sigma_{m}(.)\}_{m}$ is i.i.d.
The random variable $\widehat{\sigma}_{m}$ can be interpreted as follows. Assume that all customers of block $Y_{m}$ are already available at time $D_{m-1}$ of departure of the last customer from the previous block. These customers then form a pile of hailstones. Start serving all customers in this pile following the induction steps of Section II, namely in the order prescribed by the precedence relations between these customers, and at any time, at the individual rate prescribed by the induction steps. Then $D_{m-1}+\widehat{\sigma}_{m}\equiv D_{m}$ is the epoch when the last element of this pile leaves.
Furthermore, since the random variables $\{Y_{m}\}$ are independent for $m\geq 0$ and identically distributed for $m\geq 1$, the same holds for $\widehat{\sigma}_{m}$: they are independent for $m\geq 0$ and identically distributed for $m\geq 1$.
@@
Under the foregoing assumptions, $\mathbb{E}\widehat{\sigma}_{m}<\infty,$ for any $m$. Moreover, there exists an absolute bound, say $K$, such that ${\mathbb{E}}\Sigma_{m}(x)\leq K$, for all $x$ and $m$. In addition, the random variables $\{\Sigma_{m}(x),m\geq 1,x\geq 0\}$ are uniformly integrable and, for any $m$, we have $\Sigma_{m}(x)\to\widehat{\sigma}_{m}$ a.s., as $x\to\infty$. Therefore, for all such $m$, ${\mathbb{E}}\Sigma_{m}(x)\to{\mathbb{E}}\widehat{\sigma}_{1}$ as well.
@@
As shown in Lemma 5, the number of customers in block $m$, $\beta_{m}$, is a geometrically distributed random variable with parameter $p_{b}$, multiplied by $2k$. Further, since the total service rate (of all customers) in a non-empty system is bounded below by a positive constant, say $\kappa$ (see Lemma 1), the total time for service of all customers in a block cannot exceed the sum of their heights, divided by $\kappa$. Note also that the random variables
@@
are i.i.d.
Then, for any $m=1,2,\ldots$,
${\mathbb{E}}\sum_{j=0}^{\beta_{m}/(2k)-1}\sum_{r=0}^{2k-1}h_{n_{m}+2kj+r}$ $={\mathbb{E}}\sum_{j=1}^{\beta_{m}/(2k)}\sum_{r=-k}^{k-1}h_{n_{m}+2kj+r},$
where on the left-hand side we have the sum of the heights of customers within one block, and on the right-hand side the first $k$ terms are replaced by the first $k$ terms from the following block. For the right-hand side, given any value of $n_{m}$, the random variable $\beta_{m}/2k$ is a stopping time for the random variables $H_{j}:=\sum_{r=-k}^{k-1}h_{n_{m}+2kj+r}$. In particular, for any $n$, the event $\{\beta_{m}\leq n\}$ does not depend on the random variable $H_{n+1}$. Therefore, by the Wald identity, for $m=1,2,\ldots$,
${\mathbb{E}}\sum_{j=0}^{\beta_{m}/(2k)-1}\sum_{r=0}^{2k-1}h_{n_{m}+2kj+r}$ $={\mathbb{E}}\sum_{j=1}^{\beta_{m}/(2k)}\sum_{r=-k}^{k-1}h_{n_{m}+2kj+r}$
$=\frac{{\mathbb{E}}\beta_{m}}{2k}{\mathbb{E}}\sum_{r=0}^{2k-1}h_{r}={\mathbb{E}}\beta_{2}{\mathbb{E}}h_{1}=:C_{1}<\infty.$
In the first block, only the distribution of the $k$ random variables $\{h_{i}\}_{i=0}^{k-1}$, differs from that of the other blocks, in the sense that these random variables have the initial (unconditional) distribution. Then, by similar arguments,
${\mathbb{E}}\sum_{j=0}^{\beta_{0}/(2k)-1}\sum_{r=0}^{2k-1}h_{2kj+r}=:C_{2}<\infty.$
Furthermore, for any $m$ and $x$, the value of $\Sigma_{m}(x)$ cannot exceed $D_{m}-D_{m-1}$ (recall that $D_{m}$ is the departure time of the last customer in block $m$) and that
@@
${\mathbb{E}}\Sigma_{m}(x)\leq{\mathbb{E}}(T_{m}-T_{m-1})+C_{1}/\kappa=2k{\mathbb{E}}t_{1}/p_{b}+C_{1}/\kappa=:K<\infty.$
Next, the uniform integrability of all $\{\Sigma_{m}(x)\}$ follows directly because they are all stochastically bounded by $Z_{m}$ and with finite expectations.
@@
$\Sigma_{m}(x)=\widehat{\sigma}_{m},\ \ \text{ for all}\ \ x\geq\widehat{T}_{m}-T_{m}.$
Indeed, given $W_{m}=x$, the first customer of block $m$ starts its service at time $T_{m}+x$. So, if $x\geq\widehat{T}_{m}-T_{m}$, then all customers of block $m$ are already in the system by time $T_{m}+x$, like in the saturated system. ∎
III-D Main Stability Results
We obtain two stability results, one for the Markov chain $\{W_{n}\}$ and one for the continuous-time process $\{W(t),t\geq 0\}$ (defined later), which admits $\{W_{n}\}$ as an embedded Markov chain.
Below, we use the following notation: $T:=T_{2}-T_{1}$, $\beta:=\beta_{1}$ (where the latter is defined in the proof of Lemma 10), and $\widehat{\sigma}:=\widehat{\sigma}_{1}$. Since $T=t_{n_{2}}-t_{n_{1}}$, ${\mathbb{E}}\beta=2k/p_{b}$$\Lambda=1/\mathbb{E}t_{1}$ and $\beta$ does not depend on $\{t_{i}\}$, we get
${\mathbb{E}}T=\frac{2k}{\Lambda p_{b}}.$ (12)
@@ III-D Main Stability Results
    If $^{\mathbb{E}}T>{\mathbb{E}}\widehat{\sigma}$ – or, equivalently,
    $\Lambda<\Lambda_{c}:=\frac{2k}{p_{b}{\mathbb{E}}\widehat{\sigma}}$, (13)
@@
    In addition, the Markov chain $\{W_{n}\}$ admits a unique stationary distribution, say $\pi$, and, for any initial state $x\in{\cal X}$, the distribution of the Markov chain converges to $\pi$ in the total variation norm.
@@
    If $^{\mathbb{E}}T<{\mathbb{E}}\widehat{\sigma}$ and if there is a constant $\gamma>0$ such that, for any $x$
    ${\mathbb{P}}(\Sigma(x,Y_{1})>T+\gamma)\geq\gamma,$ (14)
@@
    $\liminf_{n\to\infty}\frac{W_{n}}{n}\geq\varepsilon\quad\text{a.s.}$
@@
The statement of Theorem 1 continues to hold true in a more general setting, for a renewal arrival process, if the distribution of $t_{1}$ has unbounded support.
Remark 12.
In some special cases, there are closed-form expressions for $p_{b}$ and $^{\mathbb{E}}\widehat{\sigma}$. For example, for ${\cal X}=[-Q,Q]^{2}$, $S_{i}=B(x_{i},R_{i})$ where $\{R_{i}\}$ are i.i.d. exponentially distributed random variables with mean $R$, it is sufficient to define the zigzag event based on a single arrival for which the exclusion set covers $\cal X$, that is, $C_{n}=\{{\cal X}\subset S_{n}\}$. In this case, $p_{b}$ is simply the tail probability of an exponential random variable, and we have:
$\Lambda_{c}=\frac{\exp(\sqrt{2}Q/R)}{{\mathbb{E}}\widehat{\sigma}}$.
The tractability of $^{\mathbb{E}}\widehat{\sigma}$ depends both on the distribution of $S_{i}$ and the pathloss function $\ell$. In general, we can estimate $p_{b}$ and $^{\mathbb{E}}\widehat{\sigma}$ by Monte Carlo simulation. The details are given in Section IV.
For any time $t\geq 0$, consider an auxiliary system in which we stop arrivals from time $t$ onward and admit only customers arriving before time $t$. In this system, there is a (random) finite number of arrivals. Denote by $Z(t)$ the time at which the last customer leaves such a system and by W(t)=Z(t)tW(t)=Z(t)-t the time to empty the system after the interruption of the input process. One can see that $Z(t)$ is finite a.s. (in particular, because the sum of the service rates of customers on the ground is bounded from below) and that $W_{n}=W(T_{n})$, for $n=1,2,\ldots$.
@@
    Assume the assumptions of item (i) of the previous theorem hold. The stochastic process $W(t)$ admits a single stationary distribution and, for any initial finite configuration of customers, the distribution of $W(t)$ converges to its stationary version in the total variation norm.
@@
    $\liminf_{n\to\infty}\frac{W(t)}{t}\geq\delta\quad\text{a.s.}$
The proofs of these two theorems are given in Appendix VII.
Remark 13.
The statement of Theorem 2 continues to hold true for a renewal arrival process if the distribution of $t_{1}$ has a spread-out distribution.
@@
IV Numerical Evaluations
In this section, we provide numerical evaluations of the system performance and compare the stability threshold with relevant models. For this comparison, we recall from Subsection II-A that $\lambda=\Lambda/m(\cal X)$ and define $\lambda_{c}={\Lambda_{c}}/{m(\cal X)}$.
IV-A Estimates of the Stability Threshold
@@
In view of the expression obtained in (13), the most natural way to estimate $\lambda_{c}$ is to estimate $p_{b}$ and $\mathbb{E}\widehat{\sigma}_{1}$ by Monte Carlo simulation. Both problems can be solved by producing perfect block samples and using them as input to the induction steps of Section II. For this purpose, in these blocks, there is no need to keep the arrival epochs. The simplest way is to first sample an i.i.d. sequence of arrival loci, height, and exclusion sets until one sees a first $C$ event and to proceed sampling such data until one sees a second $C$ event. This provides a sample of a block by restriction to arrivals that start with the zag of the first $C$ event and last until the zig of the second one. This provides a sample of the number of arrivals in a block, say $\beta$. Then we take this block as input to the induction steps of Section II. This in turn provides a perfect sample of $\widehat{\sigma}_{1}$. By repeating this, one gets an estimate (together with confidence intervals) of $\mathbb{E}\beta=2k/{p_{b}}$ and of $\mathbb{E}\widehat{\sigma}$.
IV-B Examples
@@
This subsection defines two sets of basic examples, one in the continuous space setting and the other one in the discrete space setting.
IV-B1 Continuous-Space Setting
Consider the PHWG model with a state space equal to the square window ${\cal X}=[-Q,Q]^{2}$ and $m$ the Lebesgue measure on the latter. The torus version of the model is used to avoid boundary effects.